Fermi National Laboratory
 

Operation Summary for March 8, 2010 - March 10, 2010

 

Monday March 8

The day shift began with Operations monitoring the Tevatron (TeV) store 7658 (109.57E30) and the Recycler stash (307.4E10), with the Antiproton Source stacking (10.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations disabled store 7658 at 9 AM. Operations prepared the TeV A-E enclosures, and the CDF and D0 collision halls for accesses to address detector issues and the vacuum at sector C17. Operations put the TeV into HEP shot setup at 3:16 PM.

During the evening shift, Operations established store 7661 at 4:30 PM with an Average Initial Luminosity (AIL) of 354.4E30 – This store set a New Record for the highest luminosity. FESS personnel configured the MI LCW pumps for summer mode operation, which required a short interruption in beam operations. A MI RF station (MIRF2) developed an internal modulator problem that tripped the station off numerous times. Operations called in MI RF personnel to investigate and make repairs.

Tuesday March 9

The midnight shift began with Operations monitoring store 7661 (126.16E30) and the stash (218.1E10), with Pbar stacking (17.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, RF personnel continued troubleshooting MIRF2; the station quieted down until RF personnel left for the night, but then resumed tripping off. The STATES front end caused minor problems with antiproton transfers that were resolved with a reboot.

During the day shift, Operations terminated store 7661 at 8:45 AM and put the TeV into shot setup. Operations established store 7662 at 10 AM with an AIL of 332E30. MIRF2 continued to trip and caused numerous NuMI beam permit trips; RF personnel repaired the station. The returning MTest (T995) experiment took beam from 11:46 AM to 3:43 PM. RF personnel determined that a Pbar Debuncher RF station (DRF1-6) problem was in the tunnel and would be addressed during the next Pbar access.

During the evening shift, Operations reported minor downtimes due to the Pbar beam permit tripping off due to a power supply (M:VT101) going out of tolerance, a Booster RF vacuum pump (B:IPRF19), and to problems with the Linac RF Buncher, LRF3 & 4, and a Linac Klystron RF station (KRF3).

Wednesday March 10

The midnight shift began with Operations monitoring store 7662 (80.43E30) and the stash (347.4E10), with Pbar stacking (23.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7662 at 5:40 AM. LRF5 tripped off at 1:20 AM due to bad Power Amplifier vacuum; it required a significant amount of time to recover. Operators brought the station up very slowly and reestablished beam at 4 AM. Operations established beam to MTest at 4:31 AM. Operations put the TeV into shot setup at 5:55 AM and established store 7663 at 7:05 AM with an AIL of 340E30. A short disturbance of Booster 95 degree LCW flow caused the RF gallery and Tevatron Electron Lens (TEL1) to trip off. The flow recovered quickly and operators recovered all supplies.

The Plans for Wednesday

The plans for today are to stack, stash, and store. On Thursday (3/11/10), the Run Coordinator will allow about a four-hour downtime starting at 7 AM while FESS personnel bring a rebuilt chiller online at CUB. There will be no store in the TeV, but the Recycler/MI will maintain a stash. Other areas will be allowed to conduct accesses and maintenance per Run Coordinator approval.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/10/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory